# Probing Reheating Phase via Non-Helical Magnetogenesis and Secondary Gravitational Waves

### v1

@@ IV Results and Discussion:
$\log_{10}(T_{\rm re})$ (-3,1) $-0.38^{+1.06}_{-0.17}$ $22.77\pm 7.022$
$n_{\mathrm{B}}$ (0,1.5) $1.17^{+0.01}_{-0.01}$
Table 3: In the table above, we list the posterior distributions of the parameters for our model, obtained from two different runs, $R1$ and $R2$. In $R1$, we consider the prior for the magnetic spectral index $n_{\mathrm{B}}\in(0,1.0)$, which is consistent with the $\Delta N_{\rm eff}$ bound. In $R2$, we extend the prior range of $n_{\mathrm{B}}$ to $(0,1.5)$, while keeping the same prior for the other two parameters in both cases. We also include the corresponding Bayes Factor for each run in the table.
Figure 5: In the figures presented above, we display the posterior distributions of the parameters governing the Stochastic Gravitational Waves (SGWs) induced by non-helical magnetic fields, generated using the PTARcade code Mitridate et al. (2023). The diagonal elements of the corner plot show the 1D marginalized distributions for the parameters $n_{\mathrm{B}}$, $w_{re}$, and $T_{\rm re}$. In the left figure, we consider a prior for the magnetic spectral index $n_{\mathrm{B}}\in(1,1.0)$, while in the right figure, we consider an extended prior range for $n_{\mathrm{B}}\in(0,1.5)$.
Let us remind that for $w_{re}<1/3$, the GW spectrum exhibits a blue tilt for $f>f_{\rm re}$, with $\Omega_{\mathrm{gw}}h^{2}\propto f^{2n_{\mathrm{B}}-|n_{w}|}$. Where $n_{w}=2(1-3n_{w})/(1+3n_{w})\leq 2$. Hence, for $n_{\mathrm{B}}>1.0$, the GW spectrum exhibits a blue tilt at high frequency for $f>f_{\rm re}$. Consequently, it again violates $\Delta N_{\text{eff}}$ bound. Including other PTA observations, such as those from the European PTA (EPTA) (including data from the Indian PTA (InPTA)) et al (2023), PPTA Reardon et al. (2023), or CPTA Xu et al. (2023), may suggest a slightly lower spectral index. However, we find it difficult, or one needs extreme fine-tuning of the model parameters, to be able to fit the nano-Hz observation.
@@ V Conclusions
In this study, we explore the mechanism of magnetic field generation through an $f^{2}(\phi)F_{\mu\nu}F^{\mu\nu}$ coupling during inflation and analyze the effects of non-instantaneous reheating scenarios. Specifically, we examine how the electrical conductivity during the reheating phase impacts magnetic field evolution, focusing on two limiting cases: zero and infinite conductivity. Our analysis reveals that non-instantaneous reheating plays a pivotal role in determining the present-day strength of large-scale magnetic fields.
By varying the reheating parameters—namely the effective equation of state $w_{re}$ and the reheating temperature $T_{\rm re}$—we demonstrate that it is possible to satisfy current observational bounds on magnetic fields at Mpc scales, even for blue-tilted spectra with $n<-1/2$. For such couplings, we find that the maximum achievable present-day magnetic field strength at the 1 Mpc scale is $B_{0}\sim 2.6\times 10^{-18}~{}\text{G}$, assuming a radiation-like reheating phase ($w_{re}=1/3$) with the lowest allowed reheating temperature $T_{\rm re}\simeq 0.01~{}\text{GeV}$.
In contrast, for a matter-like reheating phase ($w_{re}=0$) with the same reheating temperature, the maximum attainable field strength drops significantly to $B_{0}\simeq 3.7\times 10^{-30}~{}\text{G}$. Although this scenario features the longest reheating duration, leading to a higher conversion efficiency of electric to magnetic energy, the enhanced dilution due to background expansion outweighs this benefit, resulting in a much weaker present-day field.
Furthermore, for stiffer equations of state ($w_{re}>1/3$), the maximum magnetic field strength achievable for $n<-1/2$ couplings is around $B_{0}\simeq 10^{-23}~{}\text{G}$, while remaining consistent with current constraints on the tensor-to-scalar ratio at CMB scales.
On the other hand, couplings with $n=2$ can lead to stronger magnetic fields that exceed the current lower bounds for appropriate reheating parameters. However, such models suffer from the well-known strong coupling problem, which limits their viability despite their potential for generating larger magnetic fields.
In the second part of this work, we explored the potential detectability of our magnetogenesis scenario under generalized reheating dynamics, where both the effective equation of state parameter $w_{re}$ and the reheating temperature $T_{\rm re}$ are treated as free parameters. We find that non-instantaneous reheating has a significant impact on the resulting gravitational wave (GW) spectrum.
Specifically, for matter-like reheating ($w_{re}=0$), the combined gravitational wave (GW) spectrum—comprising both primordial GWs (PGWs) and sourced GWs (SGWs)—exhibits two distinct spectral breaks. In contrast, stiff-like reheating scenarios with $w_{re}>1/3$ produce multiple spectral breaks, which provide a clear signature for distinguishing them from other cases. Remarkably, for $w_{re}>1/3$ and coupling indices $n>1/2$, the GW spectrum can feature up to five spectral breaks. On the other hand, for models with $n<-1/2$, where electric fields dominate and electrical conductivity is negligible, the spectrum shows a maximum of three spectral breaks. These distinctive features not only serve as unique identifiers for inflationary magnetogenesis models but also offer crucial information about early-universe microphysics, such as the conductivity behavior during reheating.
We also show that in these non-trivial reheating scenarios with $w_{re}\neq 1/3$, the tensor perturbations sourced by magnetic fields at CMB scales are modified by reheating dynamics. This opens the possibility of constraining reheating through CMB observations of the tensor-to-scalar ratio or, conversely, constraining the magnetogenesis model for a given reheating scenario.
Finally, we investigated whether such magnetogenesis models can be constrained by current Pulsar Timing Array (PTA) observations. While non-helical models cannot explain the current PTA signals due to limits from $\Delta\mathrm{N}_{\mathrm{eff}}$, our analysis shows that if the conductivity during reheating is indeed small, possibly due to the absence or transition of charged particles, the resulting GW signals could pass through the sensitivity curves of upcoming detectors such as LISA, BBO, and DECIGO. These signals also exhibit unique spectral features that are sensitive not only to the initial magnetic field spectrum but also to the details of post-inflationary evolution. Our findings suggest that future GW observations could simultaneously probe inflation, reheating, and magnetogenesis, offering a new window into the early universe.
@@ Appendix A Calculation of the coefficient $\mathcal{A}_{1},~{}\mathcal{A}_{2}~{}\&~{}\mathcal{A}_{3}$:
$\mathcal{P}_{T,s}^{\mathrm{re}}=\frac{2H_{{}_{\mathrm{I}}}^{4}}{M_{\rm P}^{4}}\left(\frac{k}{k_{\mathrm{e}}}\right)^{2(\delta-2)}\mathcal{C}_{{}_{\rm B/E}}^{2}(n_{{}_{\mathrm{B}/\mathrm{E}}})\left(\frac{k}{k_{\rm e}}\right)^{2n_{{}_{\mathrm{B}/\mathrm{E}}}}\mathcal{C}_{\mathrm{m,re}}^{2}(x_{re},x_{\mathrm{end}})\mathcal{F}_{n_{{}_{\mathrm{B}/\mathrm{E}}}}(k),$ (63)
@@
[…]
@@
[…]
@@ A.0.1 Spectral shape of $\Omega_{{}_{\mathrm{GW}}}$ for $k<k_{\mathrm{re}}$ and $k\gg k_{\mathrm{re}}$
$\mathcal{P}_{T,s}^{\mathrm{re}}=\frac{2H_{{}_{\mathrm{I}}}^{4}}{M_{\rm P}^{4}}\left(\frac{k}{k_{\mathrm{e}}}\right)^{2(\delta-2)}\mathcal{C}_{{}_{\rm B/E}}^{2}(n_{{}_{\mathrm{B}/\mathrm{E}}})\left(\frac{k}{k_{\rm e}}\right)^{2n_{{}_{\mathrm{B}/\mathrm{E}}}}\mathcal{C}_{\mathrm{m,re}}^{2}(x_{re},x_{\mathrm{end}})\mathcal{F}_{n_{{}_{\mathrm{B}/\mathrm{E}}}}(k),$ (70)
@@
$\lim_{k<<k_{\mathrm{re}}}\mathcal{P}_{T,s}^{\mathrm{re}}(k,\eta_{\rm re})\simeq\mathcal{A}_{1}\,\left(\frac{H_{{}_{\mathrm{I}}}}{M_{{}_{\mathrm{P}}}}\right)^{4}\left(\frac{k_{\mathrm{re}}}{k_{\rm e}}\right)^{2(\delta-2)}\left(\frac{k}{k_{\rm e}}\right)^{2n_{{}_{\mathrm{B}/\mathrm{E}}}}\,\mathcal{C}_{{}_{\rm B/E}}^{2}(n_{{}_{\mathrm{B}/\mathrm{E}}})\mathcal{F}_{n_{{}_{\mathrm{B}/\mathrm{E}}}}(k)$ (71)

### v2

@@ IV Results and Discussion:
$\log_{10}(T_{\rm re})$ (-3,1) $0.38^{+1.06}_{-0.17}$ $22.77\pm 7.022$
$n_{\mathrm{E}}$ (0,1.5) $1.17^{+0.01}_{-0.01}$
Table 2: In the table above, we list the posterior distributions of the parameters for our model, obtained from two different runs, $R1$ and $R2$. In $R1$, we consider the prior for the electric spectral index $n_{\mathrm{E}}\in(0,1.0)$, which is consistent with the $\Delta N_{\rm eff}$ bound. In $R2$, we extend the prior range of $n_{\mathrm{E}}$ to $(0,1.5)$, while keeping the same prior for the other two parameters in both cases. We also include the corresponding Bayes Factor for each run in the table.
Figure 5: In the figures presented above, we display the posterior distributions of the parameters governing the Stochastic Gravitational Waves (SGWs) induced by non-helical magnetic fields, generated using the PTARcade code Mitridate et al. (2023). The diagonal elements of the corner plot show the 1D marginalized distributions for the parameters $n_{\mathrm{E}}$, $w_{re}$, and $T_{\rm re}$. In the left figure, we consider a prior for the electric spectral index $n_{\mathrm{E}}\in(0,1.0)$, while in the right figure, we consider an extended prior range for $n_{\mathrm{E}}\in(0,1.5)$.
Let us remind that for $w_{re}<1/3$, the GW spectrum exhibits a blue tilt for $f>f_{\rm re}$, with $\Omega_{\mathrm{gw}}h^{2}\propto f^{2n_{\mathrm{E}}-|n_{w}|}$. Where $n_{w}=2(1-3n_{w})/(1+3n_{w})\leq 2$. Hence, for $n_{\mathrm{E}}>1.0$, the GW spectrum exhibits a blue tilt at high frequency for $f>f_{\rm re}$. Consequently, it again violates $\Delta N_{\text{eff}}$ bound. Including other PTA observations, such as those from the European PTA (EPTA) (including data from the Indian PTA (InPTA) et al (2023), PPTA Reardon et al. (2023), or CPTA Xu et al. (2023), may suggest a slightly lower spectral index. However, we find it difficult, or one needs extreme fine-tuning of the model parameters, to be able to fit the nano-Hz observation.
@@ V Conclusions
In this study, we have investigated the generation of primordial magnetic fields through an inflationary coupling of the form $f^{2}(\phi)F_{\mu\nu}F^{\mu\nu}$, with a particular focus on the effects of non-instantaneous reheating. We explored how the electrical conductivity during the reheating phase influences the evolution of electromagnetic fields, analyzing two limiting cases: zero and infinite conductivity. Our analysis demonstrates that the reheating phase plays a critical role in determining the present-day strength of magnetic fields on large scales.
By varying the reheating parameters, namely the effective equation of state $w_{re}$ and the reheating temperature $T_{\rm re}$, we showed that it is possible to satisfy current observational bounds on large-scale magnetic fields within the coupling range 2.1<n<0-2.1<n<0, provided the reheating scenario involves negligible electrical conductivity. This regime is also free from strong coupling and backreaction problems. For example, under matter-like reheating ($w_{re}=0$) with $T_{\rm re}\simeq 10^{-2}\,\text{GeV}$ and $n\simeq-2.1$ ($n_{\mathrm{E}}\simeq-0.2$), the present-day magnetic field strength at the 1 Mpc scale can reach $B_{0}(1\,\text{Mpc}^{-1})\simeq 4.26\times 10^{-25}\,\text{G}.$ For a radiation-like evolution with the same parameters, the strength can be as high as $B_{0}(1\,\text{Mpc}^{-1})\simeq 3.99\times 10^{-13}\,\text{G}.$ In the case of stiff-like evolution ($w_{re}=0.5$), the maximum field strength achievable for $n_{\mathrm{E}}\simeq 0.25$ and $T_{\rm re}\simeq 10^{-2}\,\text{GeV}$ is about $B_{0}(1\,\text{Mpc}^{-1})\simeq 4.97\times 10^{-15}\,\text{G}.$ These results show that for $w_{re}>1/3$, slightly blue-tilted electric spectra can produce magnetic fields strong enough to be detectable in future gravitational wave experiments.
In the second part of this work, we explored the generation of secondary gravitational waves (SGWs) from these magnetogenesis scenarios. For $w_{re}<1/3$, the SGW spectrum peaks at high frequencies and could be detectable by future GW detectors. In contrast, for $w_{re}>1/3$, the gravitational wave spectrum undergoes non-trivial evolution and may be relevant even at CMB scales, offering potential constraints via future B-mode polarization measurements. Most notably, higher $w_{re}$ values lead to multiple spectral breaks in the present-day GW spectrum, which could be observed by experiments like LISA, BBO, DECIGO, or even SKA. While radiation-like reheating allows the largest magnetic field strengths, the resulting GW signals often fall below detector sensitivity. However, higher $w_{re}$ scenarios are not only favorable for magnetogenesis but also offer greater promise for GW detection.
In summary, our results show that to successfully realize inflationary magnetogenesis in Ratra-type models, without encountering strong coupling or backreaction issues, one must consider non-trivial reheating scenarios with negligible electrical conductivity. Furthermore, some parameter ranges allow for the simultaneous generation of detectable gravitational wave signals. Our study, therefore, provides a crucial framework for connecting inflationary magnetogenesis with future GW observations and for constraining the physics of reheating.
@@ Appendix A Calculation of the coefficient $\mathcal{A}_{1},~\mathcal{A}_{2}~\&~\mathcal{A}_{3}$:
$\mathcal{P}_{T,s}^{\mathrm{re}}=\frac{2H_{{}_{\mathrm{I}}}^{4}}{M_{\rm P}^{4}}\left(\frac{k}{k_{\rm e}}\right)^{2(\delta-2)}\mathcal{C}_{{}_{\rm B/E}}^{2}(n_{{}_{\mathrm{B}/\mathrm{E}}})\left(\frac{k}{k_{\rm end}}\right)^{2n_{{}_{\mathrm{B}/\mathrm{E}}}}\mathcal{C}_{\mathrm{m,re}}^{2}(x_{re},x_{\mathrm{end}})\mathcal{F}_{n_{{}_{\mathrm{B}/\mathrm{E}}}}(k),$ (65)
@@
$\mathcal{C}_{\mathrm{m,re}}(x_{re},x_{\mathrm{end}})=\int_{x_{\mathrm{end}}}^{x_{re}}dx_{1}\,x_{1}^{-\delta}\mathcal{G}_{\textbf{k}}(x_{re},x_{1}),$ (66)
@@
$\mathcal{C}_{\mathrm{m,re}}(x,x_{1})$ $=\frac{\pi x^{l}}{2\mathrm{sin}(l\pi)}\int dx_{1}x_{1}^{1-l-\delta}\left[J_{l}(x)J_{-l}(x_{1})-J_{-l}(x)J_{l}(x_{1})\right]$
@@ A.0.1 Spectral shape of $\Omega_{{}_{\mathrm{GW}}}$ for $k<k_{\mathrm{re}}$ and $k\gg k_{\mathrm{re}}$
$\mathcal{P}_{T,s}^{\mathrm{re}}=\frac{2H_{{}_{\mathrm{I}}}^{4}}{M_{\rm P}^{4}}\left(\frac{k}{k_{\rm e}}\right)^{2(\delta-2)}\mathcal{C}_{{}_{\rm B/E}}^{2}(n_{{}_{\mathrm{B}/\mathrm{E}}})\left(\frac{k}{k_{\rm end}}\right)^{2n_{{}_{\mathrm{B}/\mathrm{E}}}}\mathcal{C}_{\mathrm{m,re}}^{2}(x_{re},x_{\mathrm{end}})\mathcal{F}_{n_{{}_{\mathrm{B}/\mathrm{E}}}}(k),$ (72)
@@
$\lim_{k<<k_{\mathrm{re}}}\mathcal{P}_{T,s}^{\mathrm{re}}(k,\eta_{\rm re})\simeq\mathcal{A}_{1}\,\left(\frac{H_{{}_{\mathrm{I}}}}{M_{{}_{\mathrm{P}}}}\right)^{4}\left(\frac{k_{\mathrm{re}}}{k_{\rm e}}\right)^{2(\delta-2)}\left(\frac{k}{k_{\rm end}}\right)^{2n_{{}_{\mathrm{B}/\mathrm{E}}}}\,\mathcal{C}_{{}_{\rm B/E}}^{2}(n_{{}_{\mathrm{B}/\mathrm{E}}})\mathcal{F}_{n_{{}_{\mathrm{B}/\mathrm{E}}}}(k)$ (73)